OpenStudy (anonymous):

PLEASE HELP !!! REALLY NEED HELP!!! For her phone service, Teresa pays a monthly fee of $23, and she pays an additional $0.07 per minute of use.The least she has been charged in a month is $101.96.

OpenStudy (anonymous):

What are the possible numbers of minutes she has used her phone in a month? Use m for the number of minutes, and solve your inequality for m.

OpenStudy (anonymous):

$101.96 - 23 = $78.69 divided by $0.07 = m m = 1128 minutes

OpenStudy (gorv):

monthly charge=23 so it has nothing to do with minute ..she use or not she has to pay this

OpenStudy (gorv):

and remaining amount is charged only for minute

OpenStudy (gorv):

so subtract monthly charge from total to get chrge for minute

OpenStudy (gorv):

minute charge =101.96-23=????

OpenStudy (gorv):

78.96

OpenStudy (gorv):

so this was charged for minute

OpenStudy (gorv):

0.07 charge for = 1 minute 78.96 will be charged for =78.96/0.07

OpenStudy (gorv):

tell me what u got ??

OpenStudy (anonymous):

i am suppose to use an inequality symbol @gorv

OpenStudy (gorv):

ok okkkk

OpenStudy (gorv):

so total charge for minute greater than >=101.96-23

OpenStudy (anonymous):

i'm here to save the day

OpenStudy (gorv):

charge >=78.96 minute used >=78.96/0.07

OpenStudy (anonymous):

whats the question

OpenStudy (gorv):

78.96/0.07=???

OpenStudy (gorv):

@babyblue123

OpenStudy (anonymous):

ok

OpenStudy (gorv):

use your calculator

OpenStudy (gorv):

to find 78.96/0.07

OpenStudy (anonymous):

i got one

OpenStudy (anonymous):

1128

OpenStudy (gorv):

it will be =1128

OpenStudy (gorv):

so minute used >=1128

OpenStudy (anonymous):

yea i know i got the awnser

OpenStudy (gorv):

so atleast used 1128 minute

OpenStudy (anonymous):

im suppose to use an iqeualilty symbol

OpenStudy (anonymous):

ok

OpenStudy (gorv):

minute used >=1128 i thnik it is inequality

OpenStudy (anonymous):

did you get my text

OpenStudy (gorv):

or minute >1128

OpenStudy (gorv):

i already used it above

OpenStudy (gorv):

now u got it ?? can i leave ??

OpenStudy (anonymous):

its suppose to be answer than inequality and than answer

OpenStudy (anonymous):